# Another look at inference after prediction

### v3

@@ 4 Related literature
Existing PB inference literature has established connections with semi-supervised inference, semi-parametric inference, missing data, and measurement error [Angelopoulos et al., 2023a, b, Miao et al., 2023, Zrnic and Candès, 2024a, b, Angelopoulos et al., 2024]. Here we draw connections between PB inference methods and modern and classical methods from the economics and statistics literature dating back to the 1960s. These connections reveal that aspects of the PB inference problem have been explored in various areas, albeit with differences in motivation, terminology, and notation. Given the close relationship between the PPI and CC methods, future research in PB inference could benefit from a more comprehensive review of these and other related approaches than what is provided here.
4.1 Seemingly unrelated regression & its use in statistical genetics
Seemingly unrelated regression (SUR) was introduced by economist Arnold Zellner in 1962 for simultaneous estimation of multiple related regression models [Zellner, 1962, Swamy and Mehta, 1975, Schmidt, 1977]. The regression models have different outcome variables, but are assumed to have correlated errors. SUR leverages the correlation of the error terms to produce more efficient estimators compared to estimating each regression independently. In the current setting, the SUR problem can be stated formally as
@@ 4.1 Seemingly unrelated regression & its use in statistical genetics
where $\mathbb{E}(\epsilon\mid{\bf{X}})=0$ and $\mathbb{E}(\epsilon^{\prime}\mid{\bf{X}})=0.$ An estimator for $\bm{\beta}^{*}$ that is based on augmenting $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ was proposed by Conniffe [1985] and is given by:
@@
Note that $\widehat{W}^{\text{\tiny{SUR}}}$ is an estimator of the scalar $W^{\text{\tiny{SUR}}}=\frac{\mathbb{E}[\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})]}{\mathbb{E}[\text{Var}(\epsilon^{\prime}\mid{\bf{X}})]}$. Similar to the PPI and CC methods, the contribution of the augmentation term to estimation increases with the predictiveness of $\widehat{Y}$ for $Y$. In the extreme case where $\widehat{Y}$ is a perfect predictor for $Y$ such that $\widehat{Y}=Y$, $\widehat{W}^{\text{\tiny{SUR}}}=1$ and $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ is simply $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$. In the homoskedastic case when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant and do not depend on ${\bf{X}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ [Chen and Chen, 2000]. Moreover, when the conditional distribution of $(\epsilon,\epsilon^{\prime})$ given ${\bf{X}}$ is bivariate normal with mean $\bm{0}$ and has constant variance-covariance, $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ (and thus $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ as well) is the maximum likelihood estimator of $\bm{\beta}^{*}$ and fully efficient under the assumed parametric model [Conniffe, 1985]. In other scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves a lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$.
The SUR estimator was recently applied for a PB inference problem inspired by genome-wide association studies conducted with population biobanks. These studies often rely on pre-trained models to impute outcomes of interest (i.e., phenotypes or traits) as they are partially missing due to the time and expense of ascertainment [McCaw et al., 2024]. The proposed method, “SynSurr”, assumes the true and predicted outcomes follow a bivariate normal distribution. The SynSurr estimator is equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ and asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant. Building on SynSurr, Miao et al. [2024] proposed the “POP-GWAS” estimator for a variety of regression models utilizing the augmentation technique. The POP-GWAS estimator is
@@
In contrast to the SUR estimator, $\hat{\bm{\beta}}^{\text{\tiny{POP}}}$ uses $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ instead of $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ for the augmentation, and the weight, $\widehat{W}^{\text{\tiny{POP}}}$ is not generally a consistent estimator of $W^{\text{SUR}}$. $\hat{\bm{\beta}}^{\text{\tiny{POP}}}$ is therefore less efficient than $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ when the residuals are bivariate normal with constant variance-covariance and less efficient than $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ in general.
@@ 4.2 Augmented estimation for quantifying a treatment effect
Outside of statistical genetics, the augmented estimation approach used by the PPI and CC methods has been widely used for treatment effect estimation in the statistics and biostatistics literature (e.g., Leon et al. [2003], Lu and Tsiatis [2008], Tsiatis et al. [2008], Gilbert et al. [2009], Zhang and Gilbert [2010], Tian et al. [2012], Yang and Ding [2020]). Many of these approaches build on methods from the survey sampling literature (e.g., Cassel et al. [1976], Särndal et al. [2003]).
For example, Tsiatis et al. [2008] consider covariate adjusted analyses for comparing two treatments in randomized clinical trials. Letting $X$ be an indicator for randomization to the treated or control arms, the focus of inference is on $\beta$ in the model
@@
in order to estimate the treatment effect $\Delta=\mathbb{E}(Y\mid X=1)-\mathbb{E}(Y\mid X=0)$. A standard estimator of $\Delta$ is $\widehat{\Delta}=\overline{Y}^{1}-\overline{Y}^{0}$, where $\overline{Y}^{1}=n_{1}^{-1}\sum_{i=1}^{n}Y_{i}X_{i}$ and $\overline{Y}^{0}=n_{0}^{-1}\sum_{i=1}^{n}Y_{i}(1-X_{i})$ are the empirical means of $Y$ in the treated and control groups of sizes $n_{1}$ and $n_{0}$, respectively. To utilize information in baseline covariates, ${\bf{Z}}$, Tsiatis et al. [2008] propose an augmented estimator of the treatment effect given by
@@
where $\bar{X}=n^{-1}\sum_{i=1}^{n}X_{i}$ and $h^{(k)}({\bf{Z}})$ are arbitrary scalar functions of ${\bf{Z}}$ for $k=0,1$. As $X$ and ${\bf{Z}}$ are independent due to randomization, the augmentation term in 2 converges in probability to 0, analogously to the augmentation terms in the PPI and CC estimators. In the context of PB inference, pre-trained models could be used for predicting $Y$ on the basis of ${\bf{Z}}$ in the augmentation term.
4.3 Improving precision of Monte Carlo simulations with control variates
The augmentation technique has also been utilized to improve the efficiency of Monte Carlo estimation. Specifically, Rothery [1982] considers using a control variate, a statistic whose properties are known, to improve precision when estimating power via simulation. For exposition, suppose interest lies in estimating the power of a test $T$ of $H_{0}:\theta=0$ for some scalar parameter $\theta$. Let $S$ denote an existing test of the same hypothesis whose power $\mu_{Z}=\mathbb{E}_{\theta}(Z)$ is known. Define $Y$ as an indicator that test $T$ rejects and $Z$ as an indicator that test $S$ rejects. For a set of $n$ simulation replicates, the baseline estimator of power for $T$ is simply $\overline{Y}=n^{-1}\sum_{i=1}^{n}Y_{i}$ while a control variate augmented estimator is given by:
@@ 4.3 Improving precision of Monte Carlo simulations with control variates
The variance of the augmented estimator is minimized at $\lambda^{*}=\frac{\text{Cov}(Y,Z)}{\text{Var}(Z)}$ [Rothery, 1982, Davidson and MacKinnon, 1992]. The augmentation term in 3 reduces the variance of the unadjusted power estimate $\overline{Y}$ in proportion to the correlation between $Y$ and $Z$. Building on work from the survey sampling literature, a similar idea was later used for mean estimation in the semi-supervised setting by Zhang et al. [2019].
@@ 5.1 Data generation
We perform several simulation studies to compare the performance of the PPI and CC methods. In all settings, we generate two datasets: (i) a training dataset with realizations of $(Y,{\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$ to train the ML model and (ii) an independent analysis dataset for evaluating the PB inference methods containing a labeled dataset with realizations of $(Y,{\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$ and an unlabeled dataset with realizations of $({\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$. We use 10,000 replications of the analysis dataset for each scenario and a single training set as $\hat{f}$ is assumed to be fixed. The size of the training set is set to 10,000 and the size of the analysis dataset is also $n=10,000$. We let $n_{\text{\tiny{lab}}}=1000,2000,3000,4000,5000$ and $n_{\text{\tiny{unlab}}}=n-n_{\text{\tiny{lab}}}$. For both the training and analysis datasets, $({\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$ are generated from a multivariate normal distribution. Each element of $({\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$ has a mean of 0, variance of 1, and covariance of 0.4 with the remaining variables. The goal is to conduct inference on the marginal association of $Y$ and $X_{1}$ in a linear model. The parameter of interest is therefore $\beta^{*}_{1}$ defined as
$(\beta^{*}_{0},\beta^{*}_{1})=\operatorname*{arg\,min}_{\bm{\beta}}\mathbb{E}[(Y-(\beta_{0}+\beta_{1}X_{1}))^{2}].$
We consider two primary scenarios for evaluating the PB inference methods. In both scenarios, the outcome is generated as
$Y=\theta_{X_{1}}X_{1}+\theta_{\text{\tiny{INT}}}X_{1}W_{1}+{\bf{W}}^{{\sf\scriptscriptstyle{T}}}\bm{\theta}+\epsilon$
where ${\bf{W}}^{{\sf\scriptscriptstyle{T}}}=(W_{1},\dots,W_{q})$ and the error term $\epsilon$ is simulated independently from a standard normal distribution. We varied the coefficients of $\theta_{X_{1}}$, $\theta_{\text{\tiny{INT}}}$, and $\bm{\theta}$ as well as $q$ and the composition of ${\bf{Z}}$.
Scenario 1
@@
We evaluate the performance of the PPI and CC methods: $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$. We also include two comparison methods: $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, serving as the baseline, and $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$, highlighting the limitations of naive inference. We present the percent bias for each estimator. For the PPI and CC methods, we present the confidence interval (CI) length ratio, computed as the mean of the ratio of the length of the PB inference estimator’s normal-based CI relative to the length of the normal-based CI for $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ (which is simply the ratio of the asymptotic standard errors). The CI length ratio has a baseline value of 1, which indicates that the CI of the PB inference method is the same length as the CI for $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. A CI length ratio less than 1 indicates that the PB inference estimator has lower standard error than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
@@ 5.2 Results
Table 1 presents the percent bias of the various estimators across the different scenarios and labeled dataset sizes. With the exception of $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$, the methods exhibit minimal bias as they are all consistent estimators for $\beta^{*}_{1}$. The limitations of naive inference are particularly apparent in Scenarios 2a and 2b, which do not include $X_{1}$ in ML model training. $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ exhibits an average percent bias of -47% in Scenario 2a and -64% in Scenario 2b.

### v4

@@ 4 Related literature
Before evaluating the empirical properties of the CC estimator, we provide a brief overview of several closely related areas. Existing PB inference literature has established connections with semi-supervised inference, semi-parametric inference, missing data, and measurement error. 21, 29, 30, 31, 23, 32, 33 We identify additional connections between PB inference methods and modern and classical methods from the economics and statistics literature dating back to the 1960s. These connections underscore that aspects of the PB inference problem have been previously examined, albeit with differences in motivation, terminology, and notation. Given the close relationship between the PPI and CC methods, ongoing research in PB inference would benefit from a more comprehensive review of these and other related approaches.
4.1 Seemingly unrelated regression
Seemingly unrelated regression (SUR) was introduced by economist Arnold Zellner in 1962 for simultaneous estimation of multiple related regression models. 34, 35, 36 The regression models have different outcome variables, but are assumed to have correlated errors. SUR leverages the correlation between the error terms to produce more efficient estimators compared to estimating each regression separately. In the current setting, the SUR problem considers the regression models
@@ 4.1 Seemingly unrelated regression
where $\mathbb{E}(\epsilon\mid{\bf{X}})=0$ and $\mathbb{E}(\epsilon^{\prime}\mid{\bf{X}})=0.$ An estimator for $\bm{\beta}^{*}$ that is based on augmenting $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ was proposed by Conniffe 37 and is given by:
@@
Note that $\widehat{W}^{\text{\tiny{SUR}}}$ is an estimator of $W^{\text{\tiny{SUR}}}=\frac{\mathbb{E}[\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})]}{\mathbb{E}[\text{Var}(\epsilon^{\prime}\mid{\bf{X}})]}$. Similar to the PPI and CC methods, the contribution of the augmentation term to estimation increases with the predictiveness of $\widehat{Y}$ for $Y$. In the extreme case where $\widehat{Y}$ is a perfect predictor for $Y$ such that $\widehat{Y}=Y$, $\widehat{W}^{\text{\tiny{SUR}}}=1$, and $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ is simply $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$. In the homoskedastic case when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant and do not depend on ${\bf{X}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$. 22 Moreover, when the conditional distribution of $(\epsilon,\epsilon^{\prime})$ given ${\bf{X}}$ is bivariate normal with mean $\bm{0}$ and has constant variance-covariance, $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ (and thus $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ as well) is the maximum likelihood estimator of $\bm{\beta}^{*}$ and fully efficient under the assumed parametric model. 37 In other scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves a lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$.
The SUR estimator was recently applied in a PB inference problem inspired by genome-wide association studies conducted with population biobanks. These studies often rely on pre-trained models to impute outcomes of interest (i.e., phenotypes or traits) as they are partially missing due to the time and expense of ascertainment. 12 The proposed method, “SynSurr”, assumes the true and predicted outcomes follow a bivariate normal distribution. The SynSurr estimator is equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ and asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant. Building on SynSurr, Miao et al 14 proposed the “POP-GWAS” estimator for a variety of regression models utilizing the augmentation technique. The POP-GWAS estimator is
@@
In contrast to the SUR estimator, $\hat{\bm{\beta}}^{\text{\tiny{POP}}}$ uses $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ instead of $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ for the augmentation, and the weight, $\widehat{W}^{\text{\tiny{POP}}}$, is not generally a consistent estimator of $W^{\text{SUR}}$. $\hat{\bm{\beta}}^{\text{\tiny{POP}}}$ is therefore less efficient than $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ when the residuals are bivariate normal with constant variance-covariance and less efficient than $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ in general.
@@ 4.2 Augmented estimation for quantifying a treatment effect
The augmented estimation approach used by the PPI and CC methods has also been widely used for treatment effect estimation.38, 39, 40, 41, 42, 26, 43 Many of these approaches build on methods from survey sampling.44, 45 For example, Tsiatis et al 40 consider covariate adjusted analyses for comparing two treatments in randomized clinical trials. Letting $X$ be an indicator for randomization to the treated or control arms, the focus of inference is on $\beta$ in the model
@@
to estimate the treatment effect $\Delta=\mathbb{E}(Y\mid X=1)-\mathbb{E}(Y\mid X=0)$. A standard estimator of $\Delta$ is $\widehat{\Delta}=\overline{Y}^{1}-\overline{Y}^{0}$, where $\overline{Y}^{1}=n_{1}^{-1}\sum_{i=1}^{n}Y_{i}X_{i}$ and $\overline{Y}^{0}=n_{0}^{-1}\sum_{i=1}^{n}Y_{i}(1-X_{i})$ are the empirical means of $Y$ in the treated and control groups of sizes $n_{1}$ and $n_{0}$, respectively. To utilize information from baseline covariates ${\bf{Z}}$ for improving efficiency, Tsiatis et al 40 propose an augmented estimator of the treatment effect given by
@@
where $\bar{X}=n^{-1}\sum_{i=1}^{n}X_{i}$ and $h^{(k)}({\bf{Z}})$ are arbitrary scalar functions of ${\bf{Z}}$ for $k=0,1$. As $X$ and ${\bf{Z}}$ are independent due to randomization, the augmentation term in 2 converges in probability to 0, analogously to the augmentation terms in the PPI and CC estimators. The gain in efficiency similarly arises from the potential correlation between $Y$ and ${\bf{Z}}$. This technique has been applied in numerous settings, most recently for the analysis of time-to-event outcomes46 and recurrent events.47
4.3 Control variates in Monte Carlo simulations
The augmentation technique has also been utilized to improve the efficiency of Monte Carlo estimation. Specifically, Rothery 48 considers using a control variate, a statistic whose properties are known, to improve precision when estimating power via simulation. Suppose interest lies in estimating the power of a test $T$ of $H_{0}:\theta=0$ for some scalar parameter $\theta$. Let $S$ denote an existing test of the same hypothesis whose power $\mu_{Z}=\mathbb{E}_{\theta}(Z)$ is known. Define $Y$ as an indicator that test $T$ rejects and $Z$ as an indicator that test $S$ rejects. For a set of $n$ simulation replicates, the baseline estimator of power for $T$ is simply $\overline{Y}=n^{-1}\sum_{i=1}^{n}Y_{i}$ while a control variate augmented estimator is given by:
@@ 4.3 Control variates in Monte Carlo simulations
The variance of the augmented estimator is minimized at $\lambda^{*}=\frac{\text{Cov}(Y,Z)}{\text{Var}(Z)}$. 48, 49 The augmentation term in 3 reduces the variance of the unadjusted power estimate $\overline{Y}$ in proportion to the correlation between $Y$ and $Z$. Building on work from the survey sampling literature, a similar idea was later applied for semi-supervised estimation of means.50
@@ 5.1 Data generation
We next consider two primary scenarios to evaluate the performance of the PB inference methods and to demonstrate the implications of our theoretical results. In both settings, the generative model for the outcome was
$Y=\theta_{X_{1}}X_{1}+\theta_{\text{\tiny{INT}}}X_{1}W_{1}+{\bf{W}}^{\top}\bm{\theta}+\epsilon$
where ${\bf{W}}=(W_{1},\dots,W_{q})^{{\sf\scriptscriptstyle{T}}}$ and the error term $\epsilon$ was simulated independently from a standard normal distribution. We generated $({\bf{W}}^{\top},X_{1})^{{\sf\scriptscriptstyle{T}}}$ from a mean-zero multivariate normal distribution with a compound symmetric covariance structure with unit variance and pairwise correlation of 0.4. Our objective was to perform inference on the marginal effect of $X_{1}$ on $Y$ in a linear model. Specifically, we target the parameter $\beta_{1}^{*}$, defined as
$(\beta_{0}^{*},\beta_{1}^{*})=\operatorname*{arg\,min}_{\bm{\beta}}\mathbb{E}\left[(Y-(\beta_{0}+\beta_{1}X_{1}))^{2}\right].$
The predictors to train the ML model, ${\bf{Z}}$, were constructed based on ${\bf{W}}$ and $X_{1}$. The scenarios described below were obtained by varying the values of $\theta_{X_{1}}$, $\theta_{\text{\tiny{INT}}}$, and $\bm{\theta}$, as well as the dimension $q$ of ${\bf{W}}$ and the composition of the predictors ${\bf{Z}}$.
Scenario 1 (Strong and Moderate Prediction Models)
@@
In all settings, we generated two datasets from the same underlying distribution: (i) a training dataset to train the ML model and (ii) an independent analysis dataset for evaluating the PB inference methods containing a labeled dataset and an unlabeled dataset. We used 10,000 replications of the analysis dataset for each scenario and a single training set as $\hat{f}$ is assumed to be fixed. The size of the training set was 100,000 and the size of the analysis dataset was $n=10,000$. We varied the proportion of labeled data in the analysis dataset by letting $n_{\text{\tiny{lab}}}=300,600,1000$ or $2000$ in each setting.
We evaluated the performance of the PPI and CC methods: $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$. We also included three comparison methods: $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ to serve as the baseline, $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ to highlight the limitations of naive inference, and a PB inference method from a concurrent pre-print, the PoSt-Prediction Adaptive (PSPA) method.30 The PSPA method is a generalization of the PPI++ method, an extension of the original PPI estimator that adaptively minimizes the trace of the asymptotic variance-covariance of the PPI estimator. The PSPA estimator, denoted as $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$, is designed to achieve element-wise variance reduction relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ by incorporating the information from $\widehat{Y}$ directly into the estimating equation for $\bm{\beta}^{*}$. We present the percent bias for each estimator as well as the coverage probability (CP) of the 95% confidence intervals (CIs). For the four PB inference methods ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$), we also present the relative efficiency (RE) defined as the ratio of the mean squared error (MSE) of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ relative to the MSE of the PB inference estimator.
@@ 5.2 Results
Table 1 presents the percent bias of the various estimators across the different scenarios and labeled dataset sizes. With the exception of $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$, the methods exhibit minimal bias as they are all consistent estimators for the target parameter. The limitations of naive inference are particularly apparent in Scenarios 2a and 2b, which do not include $X_{1}$ in ML model training. $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ exhibits an average percent bias of -47% in Scenario 2a and -64% in Scenario 2b.
Table 2 summarizes the CPs of the 95% CIs for all methods across the various scenarios and labeled dataset sizes. As expected, $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and the four PB inference estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$) maintain coverage near the nominal level in all settings. In contrast, the 95% CIs based on $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ achieve adequate coverage only in Scenario 1a, where a strong prediction model yields negligible bias in estimating $\beta_{1}^{*}$. Coverage drops to 92% in Scenario 1b and falls to 0% in Scenarios 2a and 2b, where bias is substantial. These results underscore the risks of relying on naive inference in practice and highlight the necessity of using reliable PB inference methods.